# Stacked Intelligent Metasurface for End-to-End OFDM System

### v1

@@ II System Model
Similar to Fig. 1, Fig. 4 illustrates a DPSIM-assisted E2E OFDM system, where the TX-DPSIM and RX-DPSIM are integrated at the transmitter and receiver sides, respectively, to process the signal in the wave domain. Considering that each layer of the metasurface in DPSIM can independently control the phase for different polarization directions, the DPSIM can thus be equivalently regarded as two identical and isolated single-polarization SIMs. To facilitate understanding the model, Fig. 5 shows a schematic diagram of the HMIMO system in the two orthogonal polarization directions. We assume that $p\in\left\{0,1\right\}$ denotes two orthogonal polarization directions. Consistent with the assumptions in Section II-A, each UE integrates an identical DPSIM device, and each metasurface layer constituting the DPSIM is perfectly identical. Moreover, $A^{\mathrm{dp},\mathrm{t}}=A^{\mathrm{dp},\mathrm{t}^{\mathrm{x}}}\times A^{\mathrm{dp},\mathrm{t}^{\mathrm{y}}}$ and $A^{\mathrm{dp},\mathrm{r}}=A^{\mathrm{dp},\mathrm{r}^{\mathrm{x}}}\times A^{\mathrm{dp},\mathrm{r}^{\mathrm{y}}}$ respectively denote the number of TX dual-polarized antennas and the number of RX dual-polarized antennas per user. We further assume that TX-DPSIM comprises $L^{\mathrm{dp}}$ layers of metasurfaces, with $M^{\mathrm{dp}}=M^{\mathrm{dp},\mathrm{x}}\times M^{\mathrm{dp},\mathrm{y}}$ DPEM units on each layer, and one RX-SIM is equipped with $K^{\mathrm{dp}}$ layers of metasurfaces, with $N^{\mathrm{dp}}=N^{\mathrm{dp},\mathrm{x}}\times N^{\mathrm{dp},\mathrm{y}}$ DPEM units on each layer. For convenience of notation, we define the set of TX metasurfaces as $\mathcal{L}^{\mathrm{dp}}=\left\{0,1,2,\cdots,L^{\mathrm{dp}}\right\}$ and the set of RX metasurfaces for the $j$-th user as $\mathcal{K}_{j}^{\mathrm{dp}}=\left\{0,1,2,\cdots,K^{\mathrm{dp}}\right\}$, where $0\in\mathcal{L}^{\mathrm{dp}}$ and $0\in\mathcal{K}_{j}^{\mathrm{dp}}$ respectively represent the BS’s TX antenna metasurface and the $j$-th user’s RX antenna metasurface. The spacing between adjacent DPEM units in the TX-DPSIM is denoted as $d^{\mathrm{dp},\mathrm{t}}$, and the spacing in the RX-DPSIM is denoted as $d^{\mathrm{dp},\mathrm{r}}$.
The transmission coefficient of the $l$-th layer for TX-DPSIM and the $k$-th layer for the $j$-th user’s RX-DPSIM are defined as (16) and (17), respectively. Considering that the interaction between electromagnetic waves and environmental scatterers is the main mechanism for changing their polarization state [39], we assume that the polarization state of the electromagnetic waves remains unchanged during their propagation between DPSIM layers. Therefore, the overall coefficient matrices for TX-DPSIM and the $j$-th user’s RX-DPSIM are respectively given by (II-C) and (19), where $\mathbf{V}_{i}^{l},\ l\in\mathcal{L}^{\mathrm{dp}}$ and $\mathbf{U}_{i}^{k},\ k\in\mathcal{K}^{\mathrm{dp}}$ are calculated in the similar way as in Section II-A for (3) and (4).
@@
$\ \ \ \ \ \ \ \ \ \mathbf{T}_{i}^{\mathrm{dp}}=\left[\begin{matrix}\mathbf{\Phi}_{0}^{L}&0\\ 0&\mathbf{\Phi}_{1}^{L}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{V}_{i}^{L}&0\\ 0&\mathbf{V}_{i}^{L}\\ \end{matrix}\right]\cdots\left[\begin{matrix}\mathbf{\Phi}_{0}^{2}&0\\ 0&\mathbf{\Phi}_{1}^{2}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{V}_{i}^{2}&0\\ 0&\mathbf{V}_{i}^{2}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{\Phi}_{0}^{1}&0\\ 0&\mathbf{\Phi}_{1}^{1}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{V}_{i}^{1}&0\\ 0&\mathbf{V}_{i}^{1}\\ \end{matrix}\right]\in\mathbb{C}^{2M^{\mathrm{dp}}\times 2A^{\mathrm{dp},\mathrm{t}}},$ (18)
$\ \ \ \ \ \ \ \ \ \mathbf{R}_{i,j}^{\mathrm{dp}}=\left[\begin{matrix}\mathbf{U}_{i}^{1}&0\\ 0&\mathbf{U}_{i}^{1}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{\Psi}_{j,0}^{1}&0\\ 0&\mathbf{\Psi}_{j,1}^{1}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{U}_{i}^{2}&0\\ 0&\mathbf{U}_{i}^{2}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{\Psi}_{j,0}^{2}&0\\ 0&\mathbf{\Psi}_{j,1}^{2}\\ \end{matrix}\right]\cdots\left[\begin{matrix}\mathbf{U}_{i}^{K}&0\\ 0&\mathbf{U}_{i}^{K}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{\Psi}_{j,0}^{K}&0\\ 0&\mathbf{\Psi}_{j,1}^{K}\\ \end{matrix}\right]\in\mathbb{C}^{2A^{\mathrm{dp},\mathrm{r}}\times 2N^{\mathrm{dp}}}.$ (19)
@@
where $\mathbf{n}_{i,j}^{\mathrm{dp}}\in\mathbb{C}^{2A^{\mathrm{dp},\mathrm{r}}\times 1}$ is the receiver noise vector with distribution $\mathcal{C}\mathcal{N}(0,\sigma^{2}\mathbf{I}_{2A^{\mathrm{dp},\mathrm{r}}})$, $\mathbf{x}_{i}^{\mathrm{dp}}\in\mathbb{C}^{2A^{\mathrm{dp},\mathrm{t}}\times 1}$ is the signal vector satisfying $\mathbb{E}\left\{\mathbf{x}_{i}^{\mathrm{dp}}\left(\mathbf{x}_{i}^{\mathrm{dp}}\right)^{\mathrm{H}}\right\}=\mathbf{I}_{2A^{\mathrm{dp},\mathrm{t}}}$, $\mathbf{p}_{i}^{\mathrm{dp}}\in\mathbb{C}^{2A^{\mathrm{dp},\mathrm{t}}\times 1}$ denotes the vector of transmit powers, $\mathbf{G}_{i,j}^{\mathrm{dp}}$ is the channel coefficient matrix between the $L^{\mathrm{dp}}$-th layer metasurface of the TX-DPSIM and the $K^{\mathrm{dp}}$-th layer metasurface of the $j$-th user’s RX-DPSIM.
Analogous to the system described in Section II-A, the model for the DPSIM-assisted E2E OFDM system is given by
@@
II-D Dual-polarization channel model (DPSIM-assisted OFDM system channel model) .
@@ III E2E problem-solving
We define the received bit vector for a multi-user E2E system as $\hat{\mathbf{b}}=\left[\hat{\mathbf{b}}_{1},\hat{\mathbf{b}}_{2},\cdots,\hat{\mathbf{b}}_{J}\right]^{\mathrm{T}}$. Our goal is to recover the transmitted vector $\mathbf{b}$ from the received vector $\hat{\mathbf{b}}$ as accurately as possible. To this end, SIM (DPSIM) attempts to actively adjust its EM (DPEM) units on its metasurface to construct a matched channel, thereby suppressing bit error problems caused by factors such as multipath effects and frequency-selective fading in the physical channel. Specifically, the objective function is to minimize the bit error rate (BER) between $\mathbf{b}$ and $\hat{\mathbf{b}}$ , which is expressed as
$\small{\mathrm{BER}\left(\mathbf{b},\hat{\mathbf{b}}\right)=\frac{1}{N^{\mathrm{bit}}}}\sum_{b=1}^{N^{\mathrm{bit}}}{P\left\{\left[\mathbf{b}\right]_{b}\neq\left[\hat{\mathbf{b}}\right]_{b}\right\}},$ (28)
@@
$\mathrm{s}.\mathrm{t}.$ $\left\|\mathbf{p}_{i}\right\|_{1}=P^{\mathrm{t}},$ (29b)
$\mathbb{E}\left\{\mathbf{x}_{i}\mathbf{x}_{i}^{\mathrm{H}}\right\}=\mathbf{I}_{A^{\mathrm{t}}},$ (29c)
$\left(1-15\right),\left(28\right).$ (29d)
@@
$\mathrm{s}.\mathrm{t}.$ $\left\|\mathbf{p}_{i}^{\mathrm{dp}}\right\|_{1}=P^{\mathrm{t}},$ (30b)
$\mathbb{E}\left\{\mathbf{x}_{i}^{\mathrm{dp}}\left(\mathbf{x}_{i}^{\mathrm{dp}}\right)^{\mathrm{H}}\right\}=\mathbf{I}_{2A^{\mathrm{dp},\mathrm{t}}},$ (30c)
@@
Due to the strong coupling among the optimization variables and the non-convex unit-modulus constraints on the transmission coefficients of the metasurface EM (DPEM) elements, obtaining the globally optimal solutions of problems $\mathcal{P}1$ and $\mathcal{P}2$ is extremely challenging. Interestingly, we observe that the considered E2E system exhibits a structural similarity to an AE. Motivated by this observation, we propose to exploit deep learning techniques to train the system, aiming to realize flexible modulation and demodulation, while simultaneously establishing an autonomous control mechanism for the SIM (DPSIM).
@@
As illustrated in Fig. 6, the core idea of the proposed EMNN is to abstract the metasurface layer in the SIM (DPSIM) into several trainable hidden layers, while modeling the actual EM wave propagation as the forward propagation process in a deep learning network. A large number of transmit bit vectors $\mathbf{b}$ are randomly generated, and the mini-batch stochastic gradient descent (SGD) algorithm is employed to train the EMNN. Upon completion of the training, the weights of the corresponding hidden layers are extracted and mapped to the SIM (DPSIM) device, thereby enabling precise control of each EM (DPEM) unit. In summary, this approach essentially achieves an artificial intelligence (AI)-driven control mechanism. Furthermore, conventional DNNs are implemented between the transmit bit vectors and the transmitted signals, as well as between the received signals and the recovered bit vectors, to realize automatic modulation and demodulation.
@@
$\mathbf{x}\odot\mathbf{p}=\left[\left(\mathbf{x}_{1}\odot\mathbf{p}_{1}\right)^{\mathrm{T}},\cdots,\left(\mathbf{x}_{N^{c}}\odot\mathbf{p}_{N^{c}}\right)^{\mathrm{T}}\right]^{\mathrm{T}},$ (31)
$\mathbf{x}^{\mathrm{dp}}\odot\mathbf{p}^{\mathrm{dp}}=\left[\left(\mathbf{x}_{1}^{\mathrm{dp}}\odot\mathbf{p}_{1}^{\mathrm{dp}}\right)^{\mathrm{T}},\cdots,\left(\mathbf{x}_{N^{c}}^{\mathrm{dp}}\odot\mathbf{p}_{N^{c}}^{\mathrm{dp}}\right)^{\mathrm{T}}\right]^{\mathrm{T}}.$ (32)
The DNN consists of several linear hidden layers, all employing the ReLU activation function. Finally, a power control layer, implemented as a fixed computation layer, performs proportional amplitude scaling of the signal according to the given total transmit power $P^{\mathrm{t}}$.
Sub-network 2. TX-SIM (TX-DPSIM) NN. This is a specialized fully-connected neural network designed to establish a bijective relationship with the actual TX-SIM/TX-DPSIM hardware. The network is composed of alternating transmission layers and metasurface layers, both abstracted from the system model. The transmission layers can be directly computed from (3) and (4), and are implemented as non-trainable fixed computation layers. In contrast, the metasurface layers correspond one-to-one with the physical metasurfaces in the TX-SIM (TX-DPSIM) device and are implemented as trainable hidden layers.
Sub-network 3. Channel layer. The channel layer is implemented as a fixed computation layer, whose parameters are obtained from the physical environment through channel estimation, channel prediction, and other relevant techniques. Considering the time-varying nature of the channel, two types of channel computation layers are designed, namely, the statistical channel and the instantaneous channel. Moreover, transfer learning is employed to train the entire model, with the detailed training procedures presented in Section III-C.
Sub-network 4. RX-SIM (RX-DPSIM) NN. Similar to the design of the TX-SIM (TX-DPSIM) NN, this is also a specialized fully-connected neural network aimed at establishing a bijective relationship with the actual RX-SIM (RX-DPSIM) hardware. The difference lies in that this sub-network is further composed of $J$ modules, each corresponding to the RX-SIM (RX-DPSIM) device integrated at a specific UE. The design principles of the transmission layers and metasurface layers are identical to those described earlier and are therefore omitted here.
@@
BS DNN Liner layer + Relu $N^{\mathrm{bit}}N^{\mathrm{c}}$ $N^{\mathrm{bit}}N^{\mathrm{c}}$
@@
Transmission layer 11 $2M$ $4M$
Metasurface layer 11 $2M$ $4M$
Transmission layer 22 $2M$ $4M$
TX-SIM Metasurface layer 22 $2M$ $4M$
@@
Transmission layer $L$ $2M$ $4M$
Metasurface layer $L$ $2M$ $4M$
Channle Channle layer $2JN$ $4JN$
Metasurface layer $K$ $2JN$ $4JN$
Transmission layer $K$ $2JN$ $4JN$
Metasurface layer K1K-1 $2JN$ $4JN$
RX-SIM Transmission layer K1K-1 $2JN$ $4JN$
@@
UE DNN Liner layer + Relu $N^{\mathrm{bit}}$ $N^{\mathrm{bit}}$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \mathcal{L}(\mathbf{b},\hat{\mathbf{b}})=-\frac{1}{N^{\mathrm{batch}}}\sum_{\mathfrak{b}=1}^{N^{\mathrm{batch}}}{\sum_{b=1}^{N^{\mathrm{bit}}}{\left\{\left[\mathbf{b}_{\mathfrak{b}}\right]_{b}\log\left(\left[\hat{\mathbf{b}}_{\mathfrak{b}}\right]_{b}\right)+(1-\left[\mathbf{b}_{\mathfrak{b}}\right]_{b})\log\left(1-\left[\hat{\mathbf{b}}_{\mathfrak{b}}\right]_{b}\right)\right\}}}.$ (31)
Sub-network 5. UE-DNN. Similar to the BS-DNN, this is also a conventional DNN. The difference is that this sub-network is further divided into $J$ modules, each deployed at an individual UE to realize the mapping from the received signal $\mathbf{y}_{j}$ or $\mathbf{y}_{j}^{\mathrm{dp}}$ to the detected bit vector $\hat{\mathbf{b}}_{j}$. The DNN alternates between batch normalization and linear layers, with all linear layers employing the ReLU activation function. Finally, a sigmoid function is applied to normalize the output to the range $\left[0,1\right]$, followed by hard decision to recover the bit vector.
@@
To accurately recover the original bit vector in time-varying channels, the E2E system needs to be retrained periodically based on the CSI information. To accelerate the training process, we adopt a transfer learning approach. Specifically, a base model is first trained using the statistical CSI, and then the model is fine-tuned in real time according to the instantaneous CSI. The loss function used during the model training is the binary cross-entropy (BCE), which can be computed as in (31), where $N^{\mathrm{batch}}$ denotes the training batch size. Similar to (28), BCE is employed to evaluate the consistency between the input and the output. It holds that $\mathcal{L}(\mathbf{b},\hat{\mathbf{b}})\geqslant 0$ with equality if and only if $\mathbf{b}=\hat{\mathbf{b}}$. Therefore, a smaller BCE indicates more accurate bitstream recovery and thus better system performance.
We update the weights of each subnetwork in the EMNN online via the MBSGD algorithm, with the weight update formulas for BS-DNN, TX-SIM NN, RX-SIM NN, and UE-DNN given respectively as

### v2

@@ Abstract
Similar to Fig. 1, Fig. 4 illustrates a DPSIM-assisted E2E OFDM system, where the TX-DPSIM and RX-DPSIM are integrated at the transmitter and receiver sides, respectively, to process the signal in the wave domain. Considering that each layer of the metasurface in DPSIM can independently control the phase for different polarization directions, the DPSIM can thus be equivalently regarded as two identical and isolated single-polarization SIMs. To facilitate understanding the model, Fig. 5 presents a schematic of the HMIMO system with two orthogonal polarization directions, where we let $p\in\{0,1\}$ denote both polarizations. Consistent with the assumptions in Section II-A, each UE integrates an identical RX-DPSIM device, and each metasurface layer constituting the RX-DPSIM is perfectly identical. Moreover, $A^{\mathrm{dp},\mathrm{t}}=A^{\mathrm{dp},\mathrm{t}^{\mathrm{x}}}\times A^{\mathrm{dp},\mathrm{t}^{\mathrm{y}}}$ and $A^{\mathrm{dp},\mathrm{r}}=A^{\mathrm{dp},\mathrm{r}^{\mathrm{x}}}\times A^{\mathrm{dp},\mathrm{r}^{\mathrm{y}}}$ respectively denote the number of TX dual-polarized antennas and the number of RX dual-polarized antennas per UE. We further assume that TX-DPSIM comprises $L^{\mathrm{dp}}$ layers of metasurfaces, with $M^{\mathrm{dp}}=M^{\mathrm{dp},\mathrm{x}}\times M^{\mathrm{dp},\mathrm{y}}$ DPEM units on each layer, and one RX-DPSIM is equipped with $K^{\mathrm{dp}}$ layers of metasurfaces, with $N^{\mathrm{dp}}=N^{\mathrm{dp},\mathrm{x}}\times N^{\mathrm{dp},\mathrm{y}}$ DPEM units on each layer. For convenience of notation, we define the set of TX metasurfaces as $\mathcal{L}^{\mathrm{dp}}=\left\{0,1,2,\cdots,L^{\mathrm{dp}}\right\}$ and the set of RX metasurfaces for the $j$-th user as $\mathcal{K}_{j}^{\mathrm{dp}}=\left\{0,1,2,\cdots,K^{\mathrm{dp}}\right\}$, where $0\in\mathcal{L}^{\mathrm{dp}}$ and $0\in\mathcal{K}_{j}^{\mathrm{dp}}$ respectively represent the BS’s TX antenna metasurface and the $j$-th user’s RX antenna metasurface. The spacing between adjacent DPEM units in the TX-DPSIM is denoted as $d^{\mathrm{dp},\mathrm{t}}$, and the spacing in the RX-DPSIM is denoted as $d^{\mathrm{dp},\mathrm{r}}$.
The transmission coefficient of the $l$-th layer for TX-DPSIM and the $k$-th layer for the $j$-th user’s RX-DPSIM are defined as (16) and (17), respectively. Considering that the interaction between EM waves and environmental scatterers is the main mechanism for changing their polarization state [39], we assume that the polarization state of the EM waves remains unchanged during their propagation between DPSIM layers. Therefore, the overall coefficient matrices for TX-DPSIM and the $j$-th user’s RX-DPSIM are respectively given by (II-C) and (19), where $\mathbf{V}_{i}^{l},\ l\in\mathcal{L}^{\mathrm{dp}}$ and $\mathbf{U}_{i}^{k},\ k\in\mathcal{K}^{\mathrm{dp}}$ are calculated in the similar way as in Section II-A for (3) and (4).
@@
$\ \ \ \ \ \ \ \ \ \mathbf{T}_{i}^{\mathrm{dp}}=\left[\begin{matrix}\mathbf{\Phi}_{0}^{L}&\mathbf{0}\\ \mathbf{0}&\mathbf{\Phi}_{1}^{L}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{V}_{i}^{L}&\mathbf{0}\\ \mathbf{0}&\mathbf{V}_{i}^{L}\\ \end{matrix}\right]\cdots\left[\begin{matrix}\mathbf{\Phi}_{0}^{2}&\mathbf{0}\\ \mathbf{0}&\mathbf{\Phi}_{1}^{2}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{V}_{i}^{2}&\mathbf{0}\\ \mathbf{0}&\mathbf{V}_{i}^{2}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{\Phi}_{0}^{1}&\mathbf{0}\\ \mathbf{0}&\mathbf{\Phi}_{1}^{1}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{V}_{i}^{1}&\mathbf{0}\\ \mathbf{0}&\mathbf{V}_{i}^{1}\\ \end{matrix}\right]\in\mathbb{C}^{2M^{\mathrm{dp}}\times 2A^{\mathrm{dp},\mathrm{t}}},$ (18)
$\ \ \ \ \ \ \ \ \ \mathbf{R}_{i,j}^{\mathrm{dp}}=\left[\begin{matrix}\mathbf{U}_{i}^{1}&\mathbf{0}\\ \mathbf{0}&\mathbf{U}_{i}^{1}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{\Psi}_{j,0}^{1}&\mathbf{0}\\ \mathbf{0}&\mathbf{\Psi}_{j,1}^{1}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{U}_{i}^{2}&\mathbf{0}\\ \mathbf{0}&\mathbf{U}_{i}^{2}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{\Psi}_{j,0}^{2}&\mathbf{0}\\ \mathbf{0}&\mathbf{\Psi}_{j,1}^{2}\\ \end{matrix}\right]\cdots\left[\begin{matrix}\mathbf{U}_{i}^{K}&\mathbf{0}\\ \mathbf{0}&\mathbf{U}_{i}^{K}\\ \end{matrix}\right]\left[\begin{matrix}\mathbf{\Psi}_{j,0}^{K}&\mathbf{0}\\ \mathbf{0}&\mathbf{\Psi}_{j,1}^{K}\\ \end{matrix}\right]\in\mathbb{C}^{2A^{\mathrm{dp},\mathrm{r}}\times 2N^{\mathrm{dp}}}.$ (19)
@@
where $\mathbf{n}_{i,j}^{\mathrm{dp}}\in\mathbb{C}^{2A^{\mathrm{dp},\mathrm{r}}\times 1}$ is the receiver noise vector with distribution $\mathcal{C}\mathcal{N}(0,\sigma^{2}\mathbf{I}_{2A^{\mathrm{dp},\mathrm{r}}})$, $\mathbf{p}_{i}^{\mathrm{dp}}\in\mathbb{C}^{2A^{\mathrm{dp},\mathrm{t}}\times 1}$ denotes the vector of transmit powers, $\mathbf{G}_{i,j}^{\mathrm{dp}}$ is the channel coefficient matrix between the $L^{\mathrm{dp}}$-th layer metasurface of the TX-DPSIM and the $K^{\mathrm{dp}}$-th layer metasurface of the $j$-th user’s RX-DPSIM, $\mathbf{x}_{i}^{\mathrm{dp}}\in\mathbb{C}^{2A^{\mathrm{dp},\mathrm{t}}\times 1}$ is the signal vector satisfying $\mathbb{E}\left\{\mathbf{x}^{\mathrm{dp}}\left(\mathbf{x}^{\mathrm{dp}}\right)^{\mathrm{H}}\right\}=\mathbf{I}_{2A^{\mathrm{dp},\mathrm{t}}N^{\mathrm{c}}}$, with $\mathbf{x}^{\mathrm{dp}}=\left[\left(\mathbf{x}_{1}^{\mathrm{dp}}\right)^{\mathrm{T}},\left(\mathbf{x}_{2}^{\mathrm{dp}}\right)^{\mathrm{T}},\cdots,\left(\mathbf{x}_{N^{\mathrm{c}}}^{\mathrm{dp}}\right)^{\mathrm{T}}\right]^{\mathrm{T}}$.
Analogous to the system described in Section II-A, the bit transmission model of the DPSIM-assisted E2E OFDM system is given by
@@
II-D Dual-polarization channel model (DPSIM-assisted E2E OFDM system channel model) .
@@
We define the received bit vector for a multi-user E2E system as $\hat{\mathbf{b}}=\left[\hat{\mathbf{b}}_{1}^{\mathrm{T}},\hat{\mathbf{b}}_{2}^{\mathrm{T}},\cdots,\hat{\mathbf{b}}_{J}^{\mathrm{T}}\right]^{\mathrm{T}}$. Our goal is to recover the transmitted vector $\mathbf{b}$ from the received vector $\hat{\mathbf{b}}$ as accurately as possible. To this end, SIM (DPSIM) attempts to actively adjust the EM (DPEM) units on its metasurface to construct a matched channel, thereby suppressing bit error problems caused by factors such as multipath effects and frequency-selective fading. Specifically, the objective function is to minimize the bit error rate (BER) between $\mathbf{b}$ and $\hat{\mathbf{b}}$ , which is expressed as
$\small{\mathrm{BER}\left(\mathbf{b},\hat{\mathbf{b}}\right)=\frac{1}{N^{\mathrm{bit}}}}\sum_{b=1}^{N^{\mathrm{bit}}}{P\left\{\left[\mathbf{b}\right]_{b}\neq\left[\hat{\mathbf{b}}\right]_{b}\right\}}.$ (28)
@@
$\mathrm{s}.\mathrm{t}.$ $\sum_{i=1}^{N^{\mathrm{c}}}{\left\|\mathbf{p}_{i}\right\|_{1}}=P^{\mathrm{t}},$ (29b)
$\mathbb{E}\left\{\mathbf{xx}^{\mathrm{H}}\right\}=\mathbf{I}_{A^{\mathrm{t}}N^{\mathrm{c}}},$ (29c)
$\left(1-15\right),\left(28\right),$ (29d)
@@
$\mathrm{s}.\mathrm{t}.$ $\sum_{i=1}^{N^{\mathrm{c}}}{\left\|\mathbf{p}_{i}^{\mathrm{dp}}\right\|_{1}}=P^{\mathrm{t}},$ (30b)
$\mathbb{E}\left\{\mathbf{x}^{\mathrm{dp}}\left(\mathbf{x}^{\mathrm{dp}}\right)^{\mathrm{H}}\right\}=\mathbf{I}_{2A^{\mathrm{dp},\mathrm{t}}N^{\mathrm{c}}},$ (30c)
@@
Due to the strong coupling among the optimization variables and the non-convex unit-modulus constraints on the transmission coefficients of the metasurface EM (DPEM) units, obtaining the globally optimal solutions of problems $\mathcal{P}1$ and $\mathcal{P}2$ is extremely challenging. Interestingly, we observe that the considered E2E system exhibits a structural similarity to an AE. Motivated by this observation, we propose to exploit deep learning techniques to train the system, aiming to realize flexible modulation and demodulation, while simultaneously establishing an autonomous control mechanism for the SIM (DPSIM).
@@
As shown in Fig. 6, the core idea of the proposed EMNN is to abstract the metasurface layer in the SIM (DPSIM) into several trainable hidden layers, while modeling the actual EM wave propagation as the forward propagation process in a deep learning network. A large number of transmit bit vectors $\mathbf{b}$ are randomly generated, and the mini-batch stochastic gradient descent (MBSGD) algorithm is employed to train the EMNN. Upon completion of the training, the weights of the corresponding hidden layers are extracted and mapped to the SIM (DPSIM) device, thereby enabling precise control of each EM (DPEM) unit. In summary, this approach essentially achieves an artificial intelligence (AI)-driven control mechanism. Furthermore, conventional DNNs are implemented between the transmit bit vectors and the transmitted signals, as well as between the received signals and the recovered bit vectors, to realize automatic modulation and demodulation.
@@
$\mathbf{p}=\left[\left(\mathbf{p}_{1}\right)^{\mathrm{T}},\left(\mathbf{p}_{2}\right)^{\mathrm{T}},\cdots,\left(\mathbf{p}_{N^{\mathrm{c}}}\right)^{\mathrm{T}}\right]^{\mathrm{T}},$ (31)
$\mathbf{p}^{\mathrm{dp}}=\left[\left(\mathbf{p}_{1}^{\mathrm{dp}}\right)^{\mathrm{T}},\left(\mathbf{p}_{2}^{\mathrm{dp}}\right)^{\mathrm{T}},\cdots,\left(\mathbf{p}_{N^{\mathrm{c}}}^{\mathrm{dp}}\right)^{\mathrm{T}}\right]^{\mathrm{T}}.$ (32)
The DNN consists of several linear hidden layers, all employing the ReLU activation function. Moreover, the power control layer is a fixed computational layer that scales the signal amplitude proportionally according to the given total transmit power $P^{\mathrm{t}}$.
Sub-network 2. TX-SIM (TX-DPSIM) NN. This is a specialized fully-connected neural network designed to establish a bijective relationship with the actual TX-SIM (TX-DPSIM) hardware. The network is composed of alternating transmission layers and metasurface layers, both abstracted from the system model. The transmission layers can be directly computed from (3) and (4), and are implemented as non-trainable fixed computation layers. In contrast, the metasurface layers correspond one-to-one with the physical metasurfaces in the TX-SIM (TX-DPSIM) device and are implemented as trainable hidden layers.
Channel layer. The channel layer is a fixed computational layer obtained from the real world via techniques such as channel estimation or channel prediction. Considering the time-varying characteristics of the channel, we design two types of channel computation layers: a statistical channel and an instantaneous channel. Then, we employ transfer learning to train the entire model, and the specific training procedure will be detailed in Section III-C.
Sub-network 3. RX-SIM (RX-DPSIM) NN. Similar to the design of the TX-SIM (TX-DPSIM) NN, this is also a specialized fully-connected neural network aimed at establishing a bijective relationship with the actual RX-SIM (RX-DPSIM) hardware. The difference lies in that this sub-network is further composed of $J$ modules, each corresponding to the RX-SIM (RX-DPSIM) device integrated at a specific UE. The design principles of the transmission layers and metasurface layers are identical to those described earlier and are therefore omitted here.
@@
BS-DNN Liner layer + Relu $N^{\mathrm{bit}}N^{\mathrm{c}}$ $N^{\mathrm{bit}}N^{\mathrm{c}}$
@@
Transmission layer 11 $2M$ $4M^{\mathrm{dp}}$
Metasurface layer 11 $2M$ $4M^{\mathrm{dp}}$
Transmission layer 22 $2M$ $4M^{\mathrm{dp}}$
TX-SIM Metasurface layer 22 $2M$ $4M^{\mathrm{dp}}$
@@
NN
Transmission layer $L$ $2M$ $4M^{\mathrm{dp}}$
Metasurface layer $L$ $2M$ $4M^{\mathrm{dp}}$
Channel Channel layer $2JN$ $4JN^{\mathrm{dp}}$
Metasurface layer $K$ $2JN$ $4JN^{\mathrm{dp}}$
Transmission layer $K$ $2JN$ $4JN^{\mathrm{dp}}$
Metasurface layer K1K-1 $2JN$ $4JN^{\mathrm{dp}}$
RX-SIM Transmission layer K1K-1 $2JN$ $4JN^{\mathrm{dp}}$
@@
UE-DNN Liner layer + Relu $N^{\mathrm{bit}}$ $N^{\mathrm{bit}}$
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \mathcal{L}\left(\mathbf{b},\hat{\mathbf{b}}\right)=-\frac{1}{N^{\mathrm{batch}}}\sum_{\mathfrak{b}=1}^{N^{\mathrm{batch}}}{\sum_{b=1}^{N^{\mathrm{bit}}}{\left\{\left[\mathbf{b}_{\mathfrak{b}}\right]_{b}\log\left(\left[\hat{\mathbf{b}}_{\mathfrak{b}}\right]_{b}\right)+(1-\left[\mathbf{b}_{\mathfrak{b}}\right]_{b})\log\left(1-\left[\hat{\mathbf{b}}_{\mathfrak{b}}\right]_{b}\right)\right\}}}.$ (31)
Sub-network 4. UE-DNN. Similar to the BS-DNN, this is also a conventional DNN. The difference is that this sub-network is further divided into $J$ modules, each deployed at an individual UE to realize the mapping from the received signal $\mathbf{y}_{j}$ or $\mathbf{y}_{j}^{\mathrm{dp}}$ to the detected bit vector $\hat{\mathbf{b}}_{j}$. The DNN alternates between batch normalization and linear layers, with all linear layers employing the ReLU activation function. Finally, a sigmoid function is applied to normalize the output to the range $\left(0,1\right)$, followed by hard decision to recover the bit vector.
@@
For accurately recovering the original bit vector in time-varying channels, the E2E system needs to be retrained periodically based on CSI information. To accelerate this process, we can employ transfer learning. Specifically, a base model is first trained using the statistical CSI, and then the model is fine-tuned in real time according to the instantaneous CSI. The loss function used during the model training is the binary cross-entropy (BCE), which can be computed as in (31), where $N^{\mathrm{batch}}$ denotes the training batch size. Similar to (28), BCE is employed to evaluate the consistency between the input and the output. It holds that $\mathcal{L}(\mathbf{b},\hat{\mathbf{b}})\geqslant 0$ with equality if and only if $\mathbf{b}=\hat{\mathbf{b}}$. Therefore, a smaller BCE indicates more accurate bitstream recovery and thus better system performance.
We update the weights of each sub-network in the EMNN online via the MBSGD algorithm, with the weight update formulas for BS-DNN, TX-SIM NN, RX-SIM NN, and UE-DNN given respectively as